1. Regression Analysis Example Sales Representative Copiers Sold Sales Calls 96 41 Brian Virost The sales manager of Copier SaleBased on the data collected by the sales manager in the Copier case, it appears that she would have some difficulty predicting the sales based on only sales calls – especially at some levels of sales calls. For example, if one of her salespeople made between 70 and 80 calls, would she be able to predict a certain number of sales? Would it be easier to predict a range of sales? If yes, what would that range be?  Hint: Review the Chapter 13 PPT slides #18-27 for more information.
Answer #1

Sales Representative

Sales Calls

Copier Sold

1

96

41

2

40

41

3

104

51

4

128

60

5

164

61

6

76

29

7

72

39

8

80

50

9

36

28

10

84

43

11

180

70

12

132

56

13

120

45

14

44

31

15

84

30

Total

1440

675

Output:

Regression Statistics

Multiple R

0.865

R Square

0.748

Adjusted R Square

0.728

Standard Error

6.720

Observations

15.000

ANOVA

df

SS

MS

F

Sig.

Regression

1.000

1738.890

1738.890

38.503

0.000

Residual

13.000

587.110

45.162

Total

14.000

2326.000

Coeff

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Lower 95.0%

Upper 95.0%

Intercept

19.980

4.390

4.552

0.001

10.497

29.463

10.497

29.463

X Variable

0.261

0.042

6.205

0.000

0.170

0.351

0.170

0.351

Answer

Regression Equation: Y = a + bX

No. of Copier Sold = 19.980 + 0.261 (Sales Calls)

For 100 calls:

No. of Copier Sold = 19.980 + 0.261(100) = 19.980+26.1 = 46.08 = 46

For 70 Calls:

No. of Copier Sold = 19.980 + 0.261(70) = 38.25 = 38

For 80 Calls:

No. of Copier Sold = 19.980 + 0.261(80) = 40.86 = 41